# Nonlinear spin and orbital Rashba-Edelstein effects induced by a femtosecond laser pulse: Simulations for Au(001)

### v1

@@ IV Concluding remarks
Additionally, our simulations show that the RSOC-induced ultrafast Edelstein effects are accompanied by both longitudinal and transversal currents, which offer additional transport channels. Our findings revealed significant differences in the mechanisms governing the transversal transport of spin and orbital angular momenta. These insights deepen our comprehension of angular momentum dynamics and may pave the way for novel approaches in spintronics and orbitronics at ultrafast timescales.
@@ Acknowledgements.
This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 328545488 – TRR 227, project B04. This work was supported by the EIC Pathfinder OPEN grant 101129641 “Orbital Engineering for Innovative Electronics”.
@@ Appendix A Symmetry analysis
Without an electric field (no laser) only operations that leave the Rashba term $\partial_{z}V$ invariant are considered: $\hat{1}$, $\hat{C}_{2z}$, $\hat{m}_{yz}$, $\hat{m}_{zx}$. As a result, both spin polarization and the currents vanish.
With electric field, considering only operations that leave the Rashba term and the electric field invariant ($\hat{1}$, $\hat{m}_{zx}$) yields that $s_{y}$ is nonzero: this is the Rashba-Edelstein effect (without Rashba contribution to the spin-orbit interaction, there is no Rashba-Edelstein effect because of the then allowed additional operations $\hat{C}_{2x}$ and $\hat{m}_{xy}$). Moreover, the longitudinal probability current $j_{x}$ is nonzero, but there is no transversal probability current $j_{y}$. By combining the symmetry properties of spin-polarization components and currents we conclude that there is a $y$-spin-polarized longitudinal current $j_{x}^{y}$ as well as $x$- and $z$-spin-polarized transversal currents $j_{y}^{x}$ and $j_{y}^{z}$. As explained above, the symmetry analysis implies that a $y$-polarized longitudinal orbital current as well as $x$- and $z$-polarized transversal orbital currents are allowed.
@@
$\hat{C}_{2x}$ $-\partial_{z}V$ $E$ $s_{x}^{\mathtt{l}}$ $-s_{y}^{\mathtt{l}}$ $-s_{z}^{\mathtt{l}}$ $j_{x}^{\mathtt{l}}$ $-j_{y}^{\mathtt{l}}$
$\hat{C}_{2y}$ $-\partial_{z}V$ $-E$ $-s_{x}^{\mathtt{r}}$ $s_{y}^{\mathtt{r}}$ $-s_{z}^{\mathtt{r}}$ $-j_{x}^{\mathtt{r}}$ $j_{y}^{\mathtt{r}}$
$\hat{C}_{2z}$ $\partial_{z}V$ $-E$ $-s_{x}^{\mathtt{l}}$ $-s_{y}^{\mathtt{l}}$ $s_{z}^{\mathtt{l}}$ $-j_{x}^{\mathtt{l}}$ $-j_{y}^{\mathtt{l}}$
$\hat{m}_{xy}$ $-\partial_{z}V$ $E$ $-s_{x}^{\mathtt{r}}$ $-s_{y}^{\mathtt{r}}$ $s_{z}^{\mathtt{r}}$ $j_{x}^{\mathtt{r}}$ $j_{y}^{\mathtt{r}}$
$\hat{m}_{yz}$ $\partial_{z}V$ $-E$ $s_{x}^{\mathtt{r}}$ $-s_{y}^{\mathtt{r}}$ $-s_{z}^{\mathtt{r}}$ $-j_{x}^{\mathtt{r}}$ $j_{y}^{\mathtt{r}}$
$\hat{m}_{zx}$ $\partial_{z}V$ $E$ $-s_{x}^{\mathtt{l}}$ $s_{y}^{\mathtt{l}}$ $-s_{z}^{\mathtt{l}}$ $j_{x}^{\mathtt{l}}$ $-j_{y}^{\mathtt{l}}$
@@
For the nanoribbon with electric field, we obtain $s_{\mu}^{\mathtt{l}}=-s_{\mu}^{\mathtt{r}}$ for $\mu=x,z$ and the Rashba-Edelstein effect deduced for the 2D sample without edges. The longitudinal current is nonzero in the entire ribbon, but for the transversal one $s_{y}^{\mathtt{l}}=-s_{y}^{\mathtt{r}}$ holds. Similarly, there is a longitudinal $y$-polarized current $j_{x}^{y}$ as well as $x$- and $z$-polarized ones with opposite sign of the spin polarization in opposite halves ($j_{x}^{\mu\mathtt{l}}=-j_{x}^{\mu\mathtt{r}}$, $\mu=x,z$). The latter relation also holds for the spin-polarized transversal current, ($j_{y}^{\mu\mathtt{l}}=-j_{y}^{\mu\mathtt{r}}$, $\mu=x,z$). $j_{y}^{y\mathtt{\eta}}$, $\eta=\mathtt{l},\mathtt{r}$, vanishes.
Figure 5: Laser-induced currents in a Au(001) sample without Rashba spin-orbit coupling. Panel (a): longitudinal probability current $\braket{j_{x}}(t)$ between next-nearest neighbor sites in $x$-direction. Panels (b) and (c): transversal $z$-polarized spin and orbital angular momentum currents in $y$-direction $\braket{j_{y}^{s_{z}}}(t)$ (b) and $\braket{j_{y}^{l_{z}}}(t)$ (c), respectively. Vertical lines indicate maxima of the laser’s electric field.
@@ Appendix B Rashba-type spin-orbit coupling in real-space tight-binding form
$\hat{H}_{\mathrm{so}}$ $=\frac{1}{4c^{2}}\boldsymbol{\sigma}\cdot(\nabla V\times\hat{\boldsymbol{p}})$ (6)
(in Hartree atomic units; $c$ speed of light, $\boldsymbol{\sigma}$ vector of Pauli matrices, $V(\boldsymbol{r})$ potential, $\hat{\boldsymbol{p}}=-\mathrm{i}\nabla$ momentum operator) describes, when considering a homogeneous potential gradient in $z$-direction, RSOC [1]
@@
$\braket{\Psi}{\hat{\boldsymbol{p}}}{\Psi}$ $\approx-\mathrm{i}\sum_{\alpha;k,l}c^{\star}_{\alpha k}c_{\alpha l}\frac{\boldsymbol{d}}{d^{2}}\left(\delta_{\boldsymbol{R}_{k},\boldsymbol{R}_{l}-\boldsymbol{d}}-\delta_{\boldsymbol{R}_{k},\boldsymbol{R}_{l}}\right)$ (12)
of the momentum operator. Since the site-diagonal terms are already treated within the atomic $\boldsymbol{L}\cdot\boldsymbol{S}$ contribution to spin-orbit coupling in the Hamiltonian $\hat{H}_{0}$ (cf. Ref. 33), we are left with the site-nondiagonal contributions. The above equation thus reduces to
$\braket{\Psi}{\hat{\boldsymbol{p}}}{\Psi}$ $\approx-\mathrm{i}\sum_{\alpha;k\not=l}c^{\star}_{\alpha k}c_{\alpha l}\frac{\boldsymbol{R}_{l}-\boldsymbol{R}_{k}}{|\boldsymbol{R}_{l}-\boldsymbol{R}_{k}|^{2}}$ (13)
@@
$\braket{\Psi}{\hat{H}_{\mathrm{so}}^{\mathrm{R}}}{\Psi}$ $\approx\sum_{\alpha,\beta;k\not=l}c^{\star}_{\beta k}t_{\beta k,\alpha l}^{\mathrm{R}}c_{\alpha l}$ (14)
@@
A Rashba parameter is usually obtained from the parabolic dispersion relation $E(\boldsymbol{k})$ of a pair of Rashba-split states [4] and hence is valid only for that state pair. In the present paper, the set of Rashba parameters comprised in $\boldsymbol{\lambda}_{\alpha}^{\mathrm{R}}$ depends on the orbital-type $\alpha=\mathrm{s},\mathrm{p}_{x},\mathrm{p}_{y}$ etc. It could also depend on the pair of sites $k$ and $l$.
@@
($k\not=l$), as found in Eq. (7) (effects of in-plane potential gradients present at surfaces with three-fold rotational symmetry has been studied in Ref. 50). Moreover, ignoring the orbital dependence, the vector
@@
lies within the $xy$-plane and is perpendicular to the distance vector between the two sites. It allows to rewrite equation (16) as $t_{kl}^{\mathrm{R}}=\mathrm{i}\,\boldsymbol{\sigma}\cdot\boldsymbol{n}_{kl}$, which is utilized in model Hamiltonians [51, 52, 53, 54, 55].
@@ Appendix C Laser-induced currents without Rashba spin-orbit coupling
Concerning the laser-induced longitudinal currents, there is only a longitudinal probability current, but neither a SAM nor an OAM current. Moreover, both the transversal $x$-polarized SAM and OAM currents vanish without RSOC, and only the $z$-polarized ones remain. This is in agreement to our previous work on the ultrafast orbital Hall effect in metallic nanoribbons [16].
@@
Figure 6: Frequency analysis. (a) Fast Fourier transformations of the laser pulse and the induced $y$-components of the spin and the orbital angular momentum. (b) As (a), but for the longitudinal probability, spin, and orbital angular momentum currents flowing in $x$-direction. (c) As (b), but for the respective transversal currents flowing in $y$-direction. Gray vertical lines indicate multiples of the laser frequency.
If not only RSOC but also the atomic spin-orbit coupling is neglected, the $z$-polarized transversal SAM current also vanishes, whereas the $z$-polarized transversal OAM current within the ACA is still nonzero. These findings are in agreement with literature; Go et al. demonstrated that within the ACA, hybridization of specific orbitals gives rise to the orbital Hall effect, which is allowed in nonmagnetic centrosymmetric systems even in the absence of (atomic) spin-orbit interaction [38, 39]. Yet, the latter is required for the spin Hall effect in such systems.
@@ Appendix D Frequency analysis
The respective longitudinal $y$-polarized SAM and OAM currents, $\braket{j_{x}^{s_{y}}}$ and $\braket{j_{x}^{l_{y}}}$, have maxima concentrated at even multiples of the laser frequency [Figure 6(b); 00, $2\omega$, $4\omega$, …]. This frequency doubling with respect to the laser-induced SAM and OAM components is explained in a two-current model as follows. The sign of the laser-induced angular momentum is directly related to the orientation of the laser’s electric field, giving an oscillation with $\omega$. The longitudinal current also oscillates with $\omega$ and transports angular momentum in $+x$ direction in one half-period of the laser pulse and the opposite angular momentum along $-x$ in the next half-period. Since the transport of angular momentum in one direction is equivalent to the transport of the opposite angular momentum in the opposite direction, the SAM and OAM currents oscillate with $2\omega$. Such an effect has also been seen in the laser-induced ultrafast orbital Hall effect [16]. The zero-frequency features indicate sustained, small contributions.
Concerning the transversal SAM and OAM currents, we find that the $x$-polarized ones, $\braket{j_{y}^{s_{x}}}$ and $\braket{j_{y}^{l_{x}}}$, show maxima at even multiples of $\omega$. Recall that these are brought about by the RSOC, as are the longitudinal currents $\braket{j_{x}^{s_{y}}}$ and $\braket{j_{x}^{l_{y}}}$.

### v2

@@ IV Concluding remarks
Additionally, our simulations show that the RSOC-induced ultrafast Edelstein effects are accompanied by both longitudinal and transversal currents, which offer additional transport channels. Our findings revealed significant differences in the mechanisms governing the transversal transport of spin and orbital angular momenta. These insights deepen our comprehension of angular momentum dynamics and may pave the way for alternative approaches in spintronics and orbitronics at ultrafast timescales.
@@ Acknowledgements.
This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation)—Project-ID 328545488—TRR 227, project B04. This work was supported by the EIC Pathfinder OPEN grant 101129641 “Orbital Engineering for Innovative Electronics”.
Data Availability Statement
The data that support the findings of this article are openly available dat .
@@ Appendix A Symmetry analysis
Without an electric field (no laser, i. e. when the system is in thermal equilibrium) only operations that leave the Rashba term $\partial_{z}V$ invariant are considered: $\hat{1}$, $\hat{C}_{2z}$, $\hat{m}_{yz}$, $\hat{m}_{zx}$. As a result, both spin polarization and the currents vanish. Note that this statement only holds for the ground state scenario before the laser pulse illuminates the sample. After the laser pulse has vanished again, the system is still in a laser-excited nonequilibrium state and not in thermal equilibrium.
With electric field (with laser, i. e. when the system is in a nonequilibrium state), considering only operations that leave the Rashba term and the electric field invariant ($\hat{1}$, $\hat{m}_{zx}$) yields that $s_{y}$ is nonzero: this is the Rashba-Edelstein effect (without Rashba contribution to the spin-orbit interaction, there is no Rashba-Edelstein effect because of the then allowed additional operations $\hat{C}_{2x}$ and $\hat{m}_{xy}$; we note in passing that the laser-induced $s_{y}$ persists after the laser pulse has vanished since there is no damping mechanism considered in the simulations). Moreover, the longitudinal probability current $j_{x}$ is nonzero, but there is no transversal probability current $j_{y}$. By combining the symmetry properties of spin-polarization components and currents we conclude that there is a $y$-spin-polarized longitudinal current $j_{x}^{y}$ as well as $x$- and $z$-spin-polarized transversal currents $j_{y}^{x}$ and $j_{y}^{z}$. As explained above, the symmetry analysis implies that a $y$-polarized longitudinal orbital current as well as $x$- and $z$-polarized transversal orbital currents are allowed.
@@
$\hat{C}_{2x}$ $\partial_{z}V$ $E$ $s_{x}^{\mathtt{l}}$ $s_{y}^{\mathtt{l}}$ $s_{z}^{\mathtt{l}}$ $j_{x}^{\mathtt{l}}$ $j_{y}^{\mathtt{l}}$
$\hat{C}_{2y}$ $\partial_{z}V$ E-E $s_{x}^{\mathtt{r}}$ $s_{y}^{\mathtt{r}}$ $s_{z}^{\mathtt{r}}$ $j_{x}^{\mathtt{r}}$ $j_{y}^{\mathtt{r}}$
$\hat{C}_{2z}$ $\partial_{z}V$ E-E $s_{x}^{\mathtt{l}}$ $s_{y}^{\mathtt{l}}$ $s_{z}^{\mathtt{l}}$ $j_{x}^{\mathtt{l}}$ $j_{y}^{\mathtt{l}}$
$\hat{m}_{xy}$ $\partial_{z}V$ $E$ $s_{x}^{\mathtt{r}}$ $s_{y}^{\mathtt{r}}$ $s_{z}^{\mathtt{r}}$ $j_{x}^{\mathtt{r}}$ $j_{y}^{\mathtt{r}}$
$\hat{m}_{yz}$ $\partial_{z}V$ E-E $s_{x}^{\mathtt{r}}$ $s_{y}^{\mathtt{r}}$ $s_{z}^{\mathtt{r}}$ $j_{x}^{\mathtt{r}}$ $j_{y}^{\mathtt{r}}$
$\hat{m}_{zx}$ $\partial_{z}V$ $E$ $s_{x}^{\mathtt{l}}$ $s_{y}^{\mathtt{l}}$ $s_{z}^{\mathtt{l}}$ $j_{x}^{\mathtt{l}}$ $j_{y}^{\mathtt{l}}$
@@
For the nanoribbon with electric field, we obtain $s_{\mu}^{\mathtt{l}}=-s_{\mu}^{\mathtt{r}}$ for $\mu=x,z$ and the Rashba-Edelstein effect deduced for the 2D sample without edges. The longitudinal probability and the $y$-polarized currents are nonzero in the entire ribbon and obey the relations $j_{x}^{\mathtt{l}}=+j_{x}^{\mathtt{r}}$ and $j_{x}^{y\mathtt{l}}=+j_{x}^{y\mathtt{r}}$, but there are longitudinal $x$- and $z$-polarized currents with opposite sign of the spin polarization in opposite halves ($j_{x}^{\mu\mathtt{l}}=-j_{x}^{\mu\mathtt{r}}$, $\mu=x,z$). Similarly, for the transversal probability and $y$-polarized current $j_{y}^{\mathtt{l}}=-j_{y}^{\mathtt{r}}$ and $j_{y}^{y\mathtt{l}}=-j_{y}^{y\mathtt{r}}$ holds. However, for the other two transversal spin-polarized currents it is $j_{y}^{\mu\mathtt{l}}=+j_{y}^{\mu\mathtt{r}}$ for $\mu=x,z$.
Figure 5: Laser-induced currents in a Au(001) sample without Rashba spin-orbit coupling. Panel (a): longitudinal probability current $\braket{j_{x}}(t)$ between next-nearest neighbor sites in $x$-direction. Panels (b) and (c): transversal $z$-polarized spin and orbital angular momentum currents in $y$-direction $\braket{j_{y}^{s_{z}}}(t)$ and $\braket{j_{y}^{l_{z}}}(t)$, respectively. Like in Figure 2, thicker lines represent the data convoluted with a Gaussian, and vertical lines indicate maxima of the laser’s electric field.
@@ Appendix B Rashba-type spin-orbit coupling in real-space tight-binding form
[…]
(in Hartree atomic units; $c$ speed of light, $\boldsymbol{\sigma}$ vector of Pauli matrices, $V(\boldsymbol{r})$ potential, $\hat{\boldsymbol{p}}=-\mathrm{i}\nabla$ momentum operator) describes, when considering a homogeneous potential gradient in $z$-direction, RSOC Bychkov and Rashba (1984)
@@
$\braket{\Psi|\hat{\boldsymbol{p}}|\Psi}$ $\approx-\mathrm{i}\sum_{\alpha;k,l}c^{\star}_{\alpha k}c_{\alpha l}\frac{\boldsymbol{d}}{d^{2}}\left(\delta_{\boldsymbol{R}_{k},\boldsymbol{R}_{l}-\boldsymbol{d}}-\delta_{\boldsymbol{R}_{k},\boldsymbol{R}_{l}}\right)$ (12)
of the momentum operator. Since the site-diagonal terms are already treated within the atomic $\boldsymbol{L}\cdot\boldsymbol{S}$ contribution to spin-orbit coupling in the Hamiltonian $\hat{H}_{0}$ (cf. Ref. Konschuh et al., 2010), we are left with the site-nondiagonal contributions. The above equation thus reduces to
$\braket{\Psi|\hat{\boldsymbol{p}}|\Psi}$ $\approx-\mathrm{i}\sum_{\alpha;k\not=l}c^{\star}_{\alpha k}c_{\alpha l}\frac{\boldsymbol{R}_{l}-\boldsymbol{R}_{k}}{|\boldsymbol{R}_{l}-\boldsymbol{R}_{k}|^{2}}$ (13)
@@
$\braket{\Psi|\hat{H}_{\mathrm{so}}^{\mathrm{R}}|\Psi}$ $\approx\sum_{\alpha,\beta;k\not=l}c^{\star}_{\beta k}t_{\beta k,\alpha l}^{\mathrm{R}}c_{\alpha l}$ (14)
@@
A Rashba parameter is usually obtained from the parabolic dispersion relation $E(\boldsymbol{k})$ of a pair of Rashba-split states Ast et al. (2007) and hence is valid only for that state pair. In the present paper, the set of Rashba parameters comprised in $\boldsymbol{\lambda}_{\alpha}^{\mathrm{R}}$ depends on the orbital-type $\alpha=\mathrm{s},\mathrm{p}_{x},\mathrm{p}_{y}$ etc. It could also depend on the pair of sites $k$ and $l$.
@@
($k\not=l$), as found in Eq. (7) (effects of in-plane potential gradients present at surfaces with three-fold rotational symmetry has been studied in Ref. Premper et al., 2007). Moreover, ignoring the orbital dependence, the vector
@@
lies within the $xy$-plane and is perpendicular to the distance vector between the two sites. It allows to rewrite equation (16) as $t_{kl}^{\mathrm{R}}=\mathrm{i}\,\boldsymbol{\sigma}\cdot\boldsymbol{n}_{kl}$, which is utilized in model Hamiltonians Chen et al. (2014); Zhang et al. (2018); Zhou et al. (2019); Busch et al. (2020, 2021).
In the context of the present study, the relevant angular momentum processes—mechanisms by which angular momentum is generated, transferred, or converted between different degrees of freedom—are the dynamics of SAM and SOC. Concerning SAM dynamics, the SEE arises from the generation of a nonequilibrium spin polarization in response to an applied electric field in systems with SOC Go et al. (2020); Johansson (2024). This involves the redistribution of SAM through interactions with external fields and scattering processes. Our model explicitly includes this mechanism by treating spin degrees of freedom coupled via SOC.
The second mechanism, SOC, provides a coupling between SAM and OAM, allowing for the interconversion of linear momentum (associated with charge currents) and SAM. This interconversion is the central angular momentum transfer mechanism underlying the Edelstein effect Johansson et al. (2021). In our framework, where the electromagnetic field is treated classically and the lattice remains rigid, SOC plays the dominant role (our model includes both atomic SOC and RSOC). In more comprehensive treatments, additional channels such as OAM of electrons or angular momentum transfer to the lattice—via phonons or rigid-body rotations—may contribute. However, such processes are often neglected in effective models focused on capturing the essential physics of spintronic effects. Our model does not incorporate these mechanisms, as they are not required for describing the SEE in its fundamental form.
@@ Appendix C Laser-induced currents without Rashba spin-orbit coupling
Concerning the laser-induced longitudinal currents, there is only a longitudinal probability current, but neither a SAM nor an OAM current. Moreover, both the transversal $x$-polarized SAM and OAM currents vanish without RSOC, and only the $z$-polarized ones remain. This is in agreement to our previous work on the ultrafast orbital Hall effect in metallic nanoribbons Busch et al. (2024).
@@
If not only RSOC but also the atomic spin-orbit coupling is neglected, the $z$-polarized transversal SAM current also vanishes, whereas the $z$-polarized transversal OAM current within the ACA is still nonzero. These findings are in agreement with literature; Go et al. demonstrated that within the ACA, hybridization of specific orbitals gives rise to the orbital Hall effect, which is allowed in nonmagnetic centrosymmetric systems even in the absence of (atomic) spin-orbit interaction Go et al. (2018, 2021). Yet, the latter is required for the spin Hall effect in such systems.
Figure 6: Frequency analysis. (a) Fast Fourier transformations of the laser pulse and the induced $y$-components of the spin and the orbital angular momentum. (b) As (a), but for the longitudinal probability, spin, and orbital angular momentum currents flowing in $x$-direction. (c) As (b), but for the respective transversal currents flowing in $y$-direction. The inset in panel (c) shows the vicinity of $2\,\omega$ and $3\,\omega$ for better visibility. Gray vertical lines indicate multiples of the laser frequency.
@@ Appendix D Frequency analysis
The respective longitudinal $y$-polarized SAM and OAM currents, $\braket{j_{x}^{s_{y}}}$ and $\braket{j_{x}^{l_{y}}}$, have maxima concentrated at even multiples of the laser frequency [Figure 6(b); 0, $2\,\omega$, $4\,\omega$, …]. This frequency doubling with respect to the laser-induced SAM and OAM components is explained in a two-current model as follows. The sign of the laser-induced angular momentum is directly related to the orientation of the laser’s electric field, giving an oscillation with $\omega$. The longitudinal current also oscillates with $\omega$ and transports angular momentum in +x+x direction in one half-period of the laser pulse and the opposite angular momentum along x-x in the next half-period. Since the transport of angular momentum in one direction is equivalent to the transport of the opposite angular momentum in the opposite direction, the SAM and OAM currents oscillate with $2\,\omega$. Such an effect has also been seen in the laser-induced ultrafast orbital Hall effect Busch et al. (2024). The zero-frequency features indicate sustained, small contributions.
Concerning the transversal SAM and OAM currents [Figure 6(c)], we find that the $x$-polarized ones, $\braket{j_{y}^{s_{x}}}$ and $\braket{j_{y}^{l_{x}}}$, show maxima at even multiples of $\omega$. Recall that these are brought about by the RSOC, as are the longitudinal currents $\braket{j_{x}^{s_{y}}}$ and $\braket{j_{x}^{l_{y}}}$.